OpenStudy (anonymous):

1. Solve for x in the interval [0 , 2 π ). (a) 2sin^2x=sin x (b) 3sin^2x+cos^2x=2 (c) 1+3cosx=cos2x (d) sin2x+sin x=0

OpenStudy (mertsj):

2sin^2x-sinx=0 sinx(2sinx-1)=0 sinx=0 or 2sinx-1=0 if sinx = 0, x = 0, pi If 2sinx-1=0, sinx = 1/2 and x = pi/6, 5pi/6

OpenStudy (anonymous):

\[2\sin^2(x)=\sin(x)\] \[2\sin^2(x)-\sin(x)=0\] \[\sin(x)(2\sin(x)-1)=0\] \[\sin(x)=0, \sin(x)=\frac{1}{2}\] \[x=0, x =\pi\] for the first one \[x=\frac{\pi}{6}, x=\frac{5\pi}{6}\] for the second

OpenStudy (anonymous):

what mertsj said

OpenStudy (anonymous):

d) \[\sin(2x)+\sin(x)=0\] \[2\sin(x)\cos(x)+\sin(x)=0\] \[\sin(x)(2\cos(x)+1)=0\] \[\sin(x)=0, \cos(x)=-\frac{1}{2}\] \[x=0, x=\frac{2\pi}{3}, x=\frac{4\pi}{3}\]

OpenStudy (mertsj):

(b) replace sin^2x with 1-cos^2x and get 3(1-cos^x)+cos^2x = 2 or 3-3cos^2x+cos^2x=2 -2cos^2=-1 cos^2x = 1/2 cos x = pos or neg sqrt2/2 x = pi/4,3pi/4, 5pi/4, 7pi/4

OpenStudy (anonymous):

mertsj am not getting the same interval with u